OpenStudy (anonymous):

Mark, twelve years old is three times as old as his sister, how old will Mark be when he is twice as old as his sister? A.12 B.15 C.16 D.19 E.21

OpenStudy (shamil98):

mark is 12, his sister is 4. 8 year difference.. 16 when she is 8. . so c.

OpenStudy (anonymous):

twelve years old is three times as old as his sister, is three times as old as his sister. How old will Mark be when he is twice as old as his sister? twelve years old= 12 is three times as old as his sister= times 3 How old will Mark be when he is twice as old. 12*3=s 1/2 s=?

OpenStudy (anonymous):

If Mark is twelve, and his age is 3 times her age, that means she is 4 yrs old. In 4 years, Mark will be 16, and the sister 8. 8x2 is 16

OpenStudy (anonymous):

actually no, hold on!

OpenStudy (anonymous):

I got 8, not 16.

OpenStudy (anonymous):

(12/3)*2=8

OpenStudy (hesan):

C. 16 is the correct answer...

OpenStudy (anonymous):

you don't need wolframalpha for this! Common!

OpenStudy (anonymous):

No, the little girl is 4, meaning in 4 years she will be 8. Further on meaning that when the girl is 8, the boy is 16. 16/2=8, the correct answer is C: 16

OpenStudy (rsadhvika):

i need wolfram for doing 2+3 :o

OpenStudy (anonymous):

yes, true, the question is asking about the boy, and my answer is for girl.

OpenStudy (anonymous):

so girl is 8 and boy is 16 where as 2 yrs ago the boy is 12 and the girl is 4 8 year difference always exists.

OpenStudy (anonymous):

No no What, 4 years ago, the girl was 4 and the boy was 12, simple error

OpenStudy (anonymous):

yes, 4 years ago, my bad....

OpenStudy (anonymous):

the answer for this one is !^

OpenStudy (hesan):

THis is just a simple question... Just see if Mark is 12 years old then his sister would be 4 years according to question.. And if Marks is 16 "(12+4)" years old then his sister must be (4+4) means 8 years old... and this is what we need so when Mark is 16 uears old hw eill be twice his sister...

OpenStudy (anonymous):

yes, we got it already, tnx...